OpenStudy (chuck187):

The cost in dollars to produce x cups of lemonade is represented by the function C(x) = 10 + 0.20x. The revenue in dollars earned by selling x cups of lemonade is represented by the function R(x) = 0.50x. What is the minimum number of whole cups of lemonade that must be sold for the revenue to exceed the cost? x= __________cups

OpenStudy (mathmate):

The minimum number of cups, x, to recuperate cost is when C(x)=R(x), or \(\large 10+0.20x=0.50x\) solve for x. The minimum number of cups to exceed the cost is x+1 if x is an integer. Otherwise, round x UP to the next integer.

OpenStudy (chuck187):

I still don't understand.......... I'm pretty bad at Algebra

OpenStudy (mathmate):

Can you solve the equation for x? \(\large 10+0.20x=0.50x\)

OpenStudy (chuck187):

Nope.....

OpenStudy (mathmate):

Here's how you would solve the equation: 1. Isolate x by transposing terms involving unknown x to the left side, and constants to the right by adding and subtracting: 10+0.2x=0.5x subtract 10 from both sides 10-10+0.2x=0.5x-10 0.2x=0.5x-10 subtract 0.5x from both sides 0.2x-0.5x=0.5x-0.5x-10 -0.3x = -10 2. Solve for x by dividing both sides by the coefficient of x. -0.3x/(-0.3) = -10/(-0.3) x=-10/(-0.3) Can you find x?